A projectile is fired in such a way that it's horizontal range is equal to three times it's maximum height. What is the angle of projection?
1
Expert's answer
2020-03-09T11:08:44-0400

The horizontal range


"R=\frac{v_0^2\sin2\theta}{g}."

The maximum height


"H=\frac{v_0^2\sin^2\theta}{2g}."

Hence


"\frac{R}{H}=\frac{2\sin2\theta}{\sin^2\theta}.""\cot\theta=\frac{R}{4H}=\frac{3}{4}.""\theta=53.1^{\circ}"